# Orbital paramagnetism without density of states enhancement in nodal-line semimetal ZrSiS

### v1

@@ Abstract
Let us here discuss each contribution in $\chi_{z}^{\text{tot}}$. Since we have $\tilde{\eta}^{\text{(i)}}=0.31$ and $\tilde{\eta}^{\text{(ii)}}=0.06$, near $T=0$, $\chi_{z}^{\text{(i)}}$ is positive, while $\chi_{z}^{\text{(ii)}}$ is small and negative as shown in Fig. 3. Their sum is $\chi_{z}^{\text{(i)}}+\chi_{z}^{\text{(ii)}}=0.26\times 10^{-4}$ emu/mol, which is the main contribution for the orbital paramagnetism near $T=0$. In the temperature range of $T<300$ K, $\chi_{z}^{\text{(i)}}+\chi_{z}^{\text{(ii)}}$ is almost constant. The temperature dependence of $\chi_{z}^{\text{tot}}$ is attributed to $\chi_{z}^{\text{(iii)}}$. At low temperatures, $\chi_{z}^{\text{(iii)}}$ is negative due to the Dirac electrons in $k_{x}$-$k_{y}$ plane but small because the chemical potential is slightly outside the gap [see Fig. 1(a) and (b)]. As the temperature increases, the diamagnetism from $\chi_{z}^{\text{(iii)}}$ grows because of the smearing as expressed by Eq. (4). This leads to negative $\chi_{z}^{\text{tot}}$ when $T>120$ K.
Note that the effect of gap is not included in the model (1). However, the gap is small ($\lesssim 0.03$ eV) compared with $\Delta(=0.8$ eV), so that the effect of the gap on $\chi_{z}$ in Fig. 3 is restricted to the small range $|\mu|/\Delta\lesssim 0.04$, which will not affect our results.
Next, we consider $\chi_{x}^{\text{tot}}=\chi_{x}^{\text{(i)}}+\chi_{x}^{\text{(ii)}}+\chi_{\rm Pauli}$, where we have assumed $\chi_{x}^{\text{(iii)}}=0$ since the magnetic field is perpendicular to the nodal line. We can see that the negative value of $\chi_{x}^{\text{tot}}$ in Fig. 4 originates from $\chi_{x}$ in Fig. 3 (b).
@@
Figure 4: Temperature dependence of $\chi_{z}^{\rm tot}$ and $\chi_{x}^{\text{\rm tot}}$. Experimental data for $\chi_{z}^{\rm exp}(\bigcirc)$ and $\chi_{x}^{\rm exp}(\square)$ [30] are also shown.
@@
Discussions.— To understand the mechanism of the interband orbital paramagnetism, which is essential for $\chi_{z}$ in ZrSiS, we introduce a simple two-band model,
$H=\alpha\frac{k_{x}^{2}-k_{y}^{2}}{2}\sigma_{0}+\alpha\frac{k_{x}^{2}+k_{y}^{2}}{2}\sigma_{z}+\Delta\sigma_{x}.$ (5)
@@
This model is obtained from the small $(k_{x},k_{y})$ expansion of Eq. (1) at $(k_{x},k_{y},k_{z})=(\pm k_{\rm R},0,0)$. As shown in the inset of Fig. 5, the upper and lower bands both have saddle points. Figure 5 shows the orbital magnetic susceptibility in the $z$ direction, $\chi_{\rm orb}$, calculated by Eq. (3). The LP (intraband) contribution $\chi_{\rm LP}$ [1, 2, 39, 68] and the interband contribution ($\chi_{\rm inter}=\chi_{\rm orb}-\chi_{\rm LP}$) are also shown in Fig. 5. Our results show that this model exhibits the orbital paramagnetism near $\mu/\Delta\sim 0$ where the ground state is insulator, which is not explained by the intraband effect. The dependence on the chemical potential is similar to that observed in Fig. 3(a), indicating that the orbital paramagnetism in Fig. 3(a) is originated from the interband effect between the two saddle points. This mechanism is in sharp contrast to that of the orbital paramagnetism due to the Van Hove singularity in the single-band 2D case [38, 39] and the flat band system [40, 41], accompanied by the divergence of the DOS. Therefore, the present result of the interband orbital paramagnetism demonstrates the new mechanism that has not been known before.
@@
Figure 5: Total orbital magnetic susceptibility $\chi_{\rm orb}$, LP contribution $\chi_{\rm LP}$, and the interband contribution $\chi_{\rm inter}\equiv\chi_{\rm orb}-\chi_{\rm LP}$ of the saddle point model Eq. (5) in units of $\chi_{1}=e^{2}/\hbar^{2}$ with $k_{\rm B}T=1/40$ with cutoff $k_{c}=4$, where we have used the unit of $\alpha=\Delta=1$. Inset: Energy dispersions for Eq. (5).
@@
Summary.— We have studied the orbital magnetism in ZrSiS on the basis of the DFT calculations and the effective model. Our results elucidate three anomalies of the orbital magnetism observed, the large orbital paramagnetism without DOS enhancement, temperature dependence, and paramagnetic-to-diamagnetic anisotropy. We have found that the orbital paramagnetism in the $C_{4}$ axial direction ($k_{z}$ direction) at low temperatures is attributed to the interband effect due to the energy fluctuation along the nodal line. This mechanism is new in that it is not accompanied by the diverging enhancement of the DOS, which suppresses Pauli paramagnetism and leads to observable large orbital paramangetism. This interband effect is understood in terms of the simple two-band effective model with negative curvature originating from the energy fluctuation, which is ubiquitous in realistic nodal-line materials. Therefore, these materials are promising platform where we can observe orbital paramagnetism. Orbital paramagnetism is one of the experimentally observable prominent features of nodal-line semimetals.
@@ Acknowledgements.
We are grateful to Y. Suzumura, M. Hayashi, Z. Hiroi, T. Osada, N. Tsuji, N. Kawashima, and H. Shinaoka for their insightful discussions. This work is supported by Grants-in-Aid for Scientific Research from the Japan Society for the Promotion of Science (No. JP22J15355, No. JP18H01162, No. JP18K03482, No. JP17H02912, No. JP21H01003, No. JP22K03447, and No. JP23H04869).

### v2

@@ Abstract
We now discuss each contribution in $\chi_{z}^{\text{tot}}$. Since we have $\tilde{\eta}^{\text{(i)}}=0.31$ and $\tilde{\eta}^{\text{(ii)}}=0.06$, near $T=0$, $\chi_{z}^{\text{(i)}}$ is positive, while $\chi_{z}^{\text{(ii)}}$ is small and negative, as shown in Fig. 3(a). Their sum is $\chi_{z}^{\text{(i)}}+\chi_{z}^{\text{(ii)}}=0.26\times 10^{-4}$ emu/mol, which is the main contribution for orbital paramagnetism near $T=0$. For temperatures below 300 K, $\chi_{z}^{\text{(i)}}+\chi_{z}^{\text{(ii)}}$ is almost constant. The temperature dependence of $\chi_{z}^{\text{tot}}$ is attributed to $\chi_{z}^{\text{(iii)}}$. At low temperatures, $\chi_{z}^{\text{(iii)}}$ is negative due to Dirac electrons in the $k_{x}$-$k_{y}$ plane but small because the chemical potential is slightly outside the gap [see Fig. 1 (d)]. As the temperature increases, the diamagnetism from $\chi_{z}^{\text{(iii)}}$ grows because of the smearing as expressed by Eq. (4). This leads to negative $\chi_{z}^{\text{tot}}$ when $T>120$ K.
Note that the effect of the gap due to SOI is neglected in the model Eq. (1) used in the above calculations. This treatment is justified, as the gap is small ($\lesssim 0.03$ eV) compared with $\Delta$ (=0.8=0.8 eV), which limits its effect on $\chi_{z}$ in Fig. 3 to the narrow range $|\mu|/\Delta\lesssim 0.04$. Explicit evaluation of the effect of gap opening is provided in SM [57]. Even in the presence of band hybridization, the residual nodal-line features continue to play a crucial role, leaving a strong imprint on the interband response and still leading to paramagnetism.
Next, we consider $\chi_{x}^{\text{tot}}=\chi_{x}^{\text{(i)}}+\chi_{x}^{\text{(ii)}}+\chi_{\rm Pauli}$, where we have assumed $\chi_{x}^{\text{(iii)}}=0$ since the magnetic field is perpendicular to the Dirac cone along the nodal line (iii). We can see that the negative value of $\chi_{x}^{\text{tot}}$ in Fig. 4 originates from $\chi_{x}$ of the nodal lines (i) and (ii) shown in Fig. 3 (b). Furthermore, we find that $\chi_{x}^{\text{(i)}}+\chi_{x}^{\text{(ii)}}$ exhibits a much weaker temperature dependence than $\chi_{z}^{\text{(iii)}}$ discussed above, although in both cases the nodal lines include segments parallel to the magnetic field. The essential difference between the nodal lines (i) and (ii), on the one hand, and (iii), on the other hand, is the energy of the Dirac points relative to $E_{F}$: (i) and (ii) cross $E_{F}$, whereas (iii) lies below $E_{F}$ [See Fig. 1 (c) and (d)]. This difference leads to the difference of the temperature dependencies between $\chi_{x}^{\rm(i)}+\chi_{x}^{\rm(ii)}$ and $\chi_{z}^{\rm(iii)}$.
@@
Figure 4: Temperature dependence of $\chi_{z}^{\rm tot}$ and $\chi_{x}^{\text{\rm tot}}$. Experimental data for $\chi_{z}^{\rm exp}(\bigcirc)$ and $\chi_{x}^{\rm exp}(\square)$ [31] are also shown. The enhancement of magnetic susceptibility $\chi_{z}$ below 10 K is owing to localized magnetic impurities [31], not the orbital effect.
The models (1) and (2), based on DFT calculations, capture the essential features of the material. Our calculations achieve quantitative agreement with the experimentally measured $\chi_{z}$ and successfully reproduce the characteristic behavior of $\chi_{x}$. These results support the validity of the circular approximation for the nodal-line geometry and confirm that the energy variation and negative curvature in the dispersion are key ingredients for the observed orbital paramagnetism. We note that this quantitative agreement can be attributed to the weak dependence of $\chi_{z}^{\rm(i)}$ and $\chi_{z}^{\rm(ii)}$ on the chemical potential $\mu$, as shown in Fig. 3(a), which leads to robustness against small variations in the omitted details. In contrast, as illustrated in Fig. 3(b), $\chi_{x}^{\rm(i)}$ and $\chi_{x}^{\rm(ii)}$ exhibit stronger $\mu$ dependence, suggesting that more realistic models for nodal line (i) and (ii) would be necessary to reproduce the experimental results with higher precision.
@@
Effective model and discussions.— To understand the mechanism underlying the interband orbital paramagnetism, which is essential for $\chi_{z}$ in ZrSiS, we introduce an effective model,
$H=\alpha\frac{k_{x}^{2}-k_{y}^{2}}{2}\sigma_{0}+\alpha\frac{k_{x}^{2}+k_{y}^{2}}{2}\sigma_{z}+d\sigma_{x},$ (5)
@@
where $\alpha$ and $d$ represent the energy scale and band hybridization, respectively. This model is obtained from a small $(k_{x},k_{y})$ expansion of Eq. (1) around $(k_{x},k_{y},k_{z})=(\pm k_{\rm R},0,k_{z0})$, for some fixed $k_{z0}\neq 0$. Figure 5 shows the orbital magnetic susceptibility in the $z$ direction, $\chi_{\rm orb}$, calculated by Eq. (3). The LP (intraband) contribution $\chi_{\rm LP}$ [1, 2, 38, 71] and the interband contribution ($\chi_{\rm inter}=\chi_{\rm orb}-\chi_{\rm LP}$) are also shown in Fig. 5. Our results show that this model exhibits orbital paramagnetism near $\mu/d\sim 0$ where the ground state is insulating, which is not explained by the intraband effect. The dependence on the chemical potential resembles that observed in Fig. 3(a), clearly showing that the orbital paramagnetism in Fig. 3(a) originates from the interband effect between the saddle points. This new mechanism stands in stark contrast to orbital paramagnetism arising from the Van Hove singularity in single-band 2D systems [37, 38] or flatband systems [40, 41], as it does not rely on divergent DOS or even the presence of Fermi surfaces. The differences between these mechanisms of orbital paramagnetism are summarized in Table 1.
Table 1: Comparison of orbital paramagnetism mechanisms across different systems
System Type Fermi surface DOS
Van Hove
singularity [37]
Intraband Necessary Divergent
Flatband [41] Interband Necessary Divergent
Nodal line (present) Interband Not required Non-divergent
@@
Figure 5: Total orbital magnetic susceptibility $\chi_{\rm orb}$, LP contribution $\chi_{\rm LP}$, and the interband contribution $\chi_{\rm inter}\equiv\chi_{\rm orb}-\chi_{\rm LP}$ of the saddle point model Eq. (5) in units of $\chi_{1}=e^{2}/\hbar^{2}$ with $k_{\rm B}T=1/40$, where we have used the unit of $\alpha=d=1$. Numerical integration has been conducted on a region $|k_{x}|\leq k_{c}$ and $|k_{y}|\leq k_{c}$ with $k_{c}=4$, showing weak cutoff dependence. Inset: Energy dispersions for Eq. (5).
@@
Summary.— We have studied the orbital magnetism in ZrSiS based on the DFT calculations and an effective model. Our results elucidate three anomalies observed in the orbital magnetism, the large orbital paramagnetism without DOS enhancement, temperature dependence, and paramagnetic-to-diamagnetic anisotropy. We have found that the orbital paramagnetism in the $C_{4}$ axial direction ($k_{z}$ direction) at low temperatures is attributed to an interband effect. This mechanism is novel in that the orbital paramagnetism is not accompanied by an enhancement of DOS. Therefore, the Pauli paramagnetism is suppressed, enabling the clear observation of orbital paramagnetism. This interband effect is understood in terms of a simple effective two-band model. In this model, the negative curvature, which arises from variation of the energy of the node points along the nodal line in the material-specific model, induces the anomalous orbital response. This variation, which persists even when a small band hybridization is present, is typical of nodal-line semimetals and difficult to achieve in conventional systems. Orbital paramagnetism is one of the most prominent features of nodal-line semimetals, and importantly, it is experimentally accessible.
@@ Acknowledgements.
We are grateful to Y. Suzumura, M. Hayashi, Z. Hiroi, T. Osada, N. Tsuji, N. Kawashima, and H. Shinaoka for their insightful discussions. This work was supported by Grants-in-Aid for Scientific Research from the Japan Society for the Promotion of Science (No. JP22J15355, No. JP18H01162, No. JP21H05191, No. JP18K03482, No. JP17H02912, No. JP21H01003, No. JP22K03447, and No. JP23H04869).